What is the energy difference between the F2 molecule and the separated atoms?

Express your answer in kilojoules per mole to three significant figures.
Answer #1

We can write the reaction:

F2(g) ----> 2F (g)

Hrxn = the energy difference between seprated atoms and F2 molecule = 2* Hf(F(g)) - Hf(F2(g))

But we know F2 is is in its natural diatomic state hence Hf(F2(g)) = 0

whereas, litrature give Hf(F(g)) = 78.99 kJ/mol

Thus Hrxn = 2*78.99 kJ/mol - 0 = 157.98 kJ/mol

So the energy difference between the F2 molecule and seprated atoms is 157.98 kJ/mol.